# Effects of particle-hole fluctuations on the superfluid transition in two-dimensional atomic Fermi gases

### v1

@@ III.3 Comparison with different results
Figure 6: Comparison of theoretical calculations for $T_{\text{BKT}}/T_{\text{F}}$ with experiment and QMC results as a function of $\ln(k_{\text{F}}a_{\text{2D}})$ throughout the BCS-BEC crossover. (a) Overlay of theoretical $T_{\text{BKT}}$ without (black solid line) and with the particle-hole channel calculated with the level 1 average (light cyan solid line), on top of the contour plot of experimentally measured quasi-condensate fractions [16]. Adapted from Ref. [18]. (b) Overlay of our $T_{\text{BKT}}$ on top of a collection of experimental data [16, 50] and various theoretical results [51, 22, 21, 23, 52]. The black solid and dashed lines represent $T_{\text{BKT}}$ calculated using our pairing fluctuation theory with and without the particle-hole channel contributions, respectively. Here the particle-hole channel effect was calculated with level 1 averaging of $\langle\chi_{\text{ph}}\rangle$. Adapted from Ref. [52].
@@
Finally, in Fig. 6, we compare our theoretical $T_{\text{BKT}}$ with available experimental data [16] and QMC results on the BKT transition in 2D Fermi gases, as a function of $\ln(k_{\text{F}}a_{\text{2D}})$ throughout the BCS-BEC crossover. Fig. 6(a) presents the measured quasi-condensate fractions $N_{q}/N$ [16], overlaid on top of which are the theory curves calculated using our pairing fluctuation theory without (black solid) and with (light cyan solid curve) the particle-hole contributions (level 1). The black solid line was previously presented in Ref. [18]. The experimental data are not dense enough to enable a successful detection of the minimum in $T_{\text{BKT}}$ near $\ln(k_{\text{F}}a_{\text{2D}})=0$, nevertheless, it does seem to suggest there is a minimum in the contour of the quasi-condensate fraction. Thus, our theory is consistent with the data, including the presence of the minimum. In Fig. 6(b), we compare our theoretical result of $T_{\text{BKT}}$ with a collection of other theories [51, 22, 21, 23], and results obtained from Quantum Monte Carlo (QMC) using a 2D lattice model [52]. Also plotted are the experimental results from Refs. [16] and [50]. In the BEC regime, our results both with (black solid) and without (black dashed line) are in quantitative agreement with both experiments. Here the particle-hole susceptibility is averaged at level 1. On the BCS side, the curves of Mulkerin et al [23] and the BCS mean-field treatment are close to our results without particle-hole, suggesting that particle-hole contributions are not included in Mulkerin et al’s calculations [23]. Our result with particle-hole contributions (black solid) are in good agreement with the QMC result in the infinite $L$ limit, except that QMC does not show a minimum in the unitary regime [52]. The results of Bauer et al [21] and Petrov et al [51] are in good agreement with QMC only in the very weak BCS regime, $\ln(k_{\text{F}}a_{\text{2D}})>2$. We also notice there is a large difference between the finite ($L=45$) and infinite lattice QMC results. It is unusual that the result of Bighin et al [22] is far above the BCS mean-field treatment in the BCS limit. It should be noted that the experimental data in the BCS regime are significantly above all theory curves, except the mean-field result. This suggests that some other factors, e.g., nonequilibrium, finite size effect, needs to be considered. Therefore, our result should not be compared with the experimental data in the BCS regime. In short, our results with the particle-hole contributions are consistent with experiment in the unitary and BEC regimes and in good agreement with QMC results in the unitary and BCS regimes. This calls for more data from future experiment.
@@
Note that since the particle-hole susceptibility in Eq. (8) is density independent in 2D, the particle-hole contributions in the weak coupling limit would have already been automatically included, if the 2D scattering length $a_{\text{2D}}$ were measured directly through experiment, e.g., by measuring the two-body binding energy in the dilute limit. Instead, $a_{\text{2D}}$ is usually calculated from the 3D scattering length as in a deeply confined pancake-shaped trap [53], thus the effect of particle-hole fluctuations should be seriously taken into account when comparing experiment and theory. In Fig. 6, the same definition of $a_{\text{2D}}$ was used in both QMC and our present work. This makes it possible to have a good agreement in the BCS regime between QMC and our result with particle-hole fluctuations included. Conversely, one can experimentally determine the particle-hole susceptibility in the BCS limit by measuring the 2D scattering length and comparing with that calculated from $a_{\text{3D}}$, and obtain
@@
where $a^{\text{exp}}_{\text{2D}}$ denotes the experimentally measured 2D scattering length. Nevertheless, away from the weak coupling limit, a nontrivial density dependence should emerge as a nonzero pairing gap develops with increasing pairing strength at a finite density.
@@ IV Conclusions
In summary, we have studied the impact of the particle-hole channel on BKT physics in Fermi gases within the context of the BCS-BEC crossover. We introduce the particle-hole channel effect by incorporating an average particle-hole susceptibility, which self-consistently includes the self-energy feedback, leading to an effective renormalization of the pairing strength. The dynamic structure of the angular-averaged particle-hole susceptibility exhibits strong dependencies on momentum and temperature, showing distinct momentum dependencies at low temperatures compared to its 3D counterpart. Furthermore, we perform averaging of the particle-hole susceptibility at two different levels, revealing important physical consequences in the crossover and BCS regimes. The particle-hole channel provides a screening of the pairing interaction and shifts the BKT transition temperature and the pairing gap curves towards the BEC regime as a function of $\ln(k_{F}a_{\text{2D}})$. Additionally, a comparison shows that the BKT transition temperature calculated with the particle-hole channel provides a better fit for the experimental data and QMC results. Future experiments with more elaborate measurements are called for to resolve various discrepancies.
Finally, it should be mentioned that, in addition to the simple averaged particle-hole susceptibility, there are a series of higher order corrections, including a higher order $T$-matrix [32], which may contribute significant modifications to the present results. In addition, the BKT criterion [24, 18] is merely based on numerical simulations [25] to provide a value of the phase space density. Despite the experimental support [17], an analytical derivation of such a criterion would be highly desirable.
@@ Appendix A Slope discontinuity in $\chi_{\text{ph}}$ across $\mu=0$ at zero $T$
Appendix B Analytical result of 2D gap equation in the BCS limit
To obtain an analytical result of the BCS gap equation for the 2D case, we define $\epsilon=\epsilon_{\mathbf{k}}/E_{\text{F}}$, $t=T/T_{\text{F}}$, and $a=\epsilon_{\text{B}}/E_{\text{F}}$. Since $\mu\approx E_{F}$, we have
$\sum_{\mathbf{k}}\left[\frac{1-2f(\xi_{\mathbf{k}})}{2\xi_{\mathbf{k}}}-\frac{1}{2\epsilon_{\mathbf{k}}+\epsilon_{\text{B}}}\right]$
$=$ $\frac{m}{2\pi}\int_{0}^{\infty}d\epsilon\left[\frac{\tanh\frac{\epsilon-1}{2t}}{2(\epsilon-1)}-\frac{1}{2\epsilon+a}\right]$
$=$ $\frac{m}{4\pi}\int_{-\frac{1}{2t}}^{\infty}dx\left[\frac{\tanh x}{x}-\frac{1}{x+\frac{1}{2t}+\frac{a}{4t}}\right]=\frac{m}{4\pi}\ln\left(\frac{2e^{2\gamma}a}{\pi^{2}t^{2}}\right).$
Here we have utilized the fact that ${1}/{2t}\gg 1$ to justify the approximation $\tanh({1}/{2t})=1$.

### v2

@@ III.3 Comparison with different results
Figure 6: Comparison of theoretical calculations for $T_{\text{BKT}}/T_{\text{F}}$ with experiment [44, 47] and QMC [23] results as a function of $\ln(k_{\text{F}}a_{\text{2D}})$ throughout the BCS-BEC crossover. (a) Overlay of theoretical $T_{\text{BKT}}$ without (black solid line, with fixed ${\cal D}_{\text{B}}=4.9$ from Ref. [49]) and with the particle-hole channel calculated with the level 1 average (light cyan solid line), on top of the contour plot of experimentally measured quasi-condensate fractions [44]. Adapted from Ref. [49]. (b) Overlay of our $T_{\text{BKT}}$ on top of a collection of experimental data [44, 47], QMC [23] and various theoretical results [39, 4, 2, 37]. The black solid and dashed lines represent $T_{\text{BKT}}$ calculated using our pairing fluctuation theory with and without the particle-hole channel contributions, respectively. Here the particle-hole channel effect was calculated with level 1 averaging of $\langle\chi_{\text{ph}}\rangle$. Adapted from Ref. [23].
@@
Finally, in Fig. 6, we compare our theoretical $T_{\text{BKT}}$ with available experimental data [44] and QMC results, as well as other theories where appropriate, on the BKT transition in 2D Fermi gases, as a function of $\ln(k_{\text{F}}a_{\text{2D}})$ throughout the BCS-BEC crossover. Figure 6(a) presents the measured quasi-condensate fractions $N_{q}/N$ [44], overlaid on top of which are the theory curves calculated using our pairing fluctuation theory without (black solid) and with (light cyan solid curve) the particle-hole contributions (level 1). The black solid line was previously presented in Ref. [49], calculated with a fixed ${\cal D}_{\text{B}}=4.9$. The experimental data are not dense enough to enable a successful detection of the minimum in $T_{\text{BKT}}$ near $\ln(k_{\text{F}}a_{\text{2D}})=0$, nevertheless, it does seem to suggest there exists a minimum in this neighborhood in the contour of the quasi-condensate fraction. Thus, our theory is consistent with the data, including the presence of the minimum. Our present curve does align better with the slightly decreasing $T_{\text{BKT}}$ and contour of the quasi-condensate fraction toward BEC. In Fig. 6(b), we compare our $T_{\text{BKT}}$ with a collection of other theories [39, 4, 2, 37], and results obtained from QMC using a 2D lattice model [23]. Also plotted are the experimental results from Refs. [44] and [47]. In the BEC regime, where particle-hole contributions become negligible, our results both with (black solid) and without (black dashed line) are in quantitative agreement with both experiments, as well as that of Petrov et al [39]. Here we choose to show only the result with level 1 average of the particle-hole susceptibility, giving the relatively small quantitative difference in $T_{\text{BKT}}$ between levels 1 and 2. There is an expected small difference between Petrov et al [39] and our present result based on experimental ${\cal D}_{\text{B}}$, due to the quantitative difference between the pair size used in Ref. [39] and the scattering length $a_{\text{2D}}$. In the BCS regime, the curves of Mulkerin et al [37] and of the BCS mean-field treatment are close to our curve without particle-hole, suggesting that particle-hole contributions are not included in Mulkerin et al [37]. Our particle-hole corrected result (black solid curve) is in good agreement with the QMC result extrapolated to the infinite $L$ limit, except that QMC does not show a minimum in the unitary regime [23]. For $\ln(k_{\text{F}}a_{\text{2D}})>2$, the QMC for infinite lattice size $L$ does not have data points, while the results of Petrov et al [39] starts to merge with our (black solid) particle-hole corrected curve. In comparison, Bauer et al [2] has a higher $T_{\text{BKT}}$ in this regime; their curve intersects with the curves of our particle-hole corrected result, Petrov et al and QMC (infinite $L$). We also notice a large difference between the finite ($L=45$) and infinite lattice QMC results. It is unusual that the result of Bighin et al [4] is far above the BCS mean-field value in the BCS limit  [^3] We note that, in Ref. [4], the expressions for the normal fluid densities $n_{n,f}$ and $n_{n,b}$ in their Eqs. (15-16) were wrong by a dimensionful factor $1/m$.. They also ignored the particle-hole contributions. It should be noted that the experimental data in the BCS regime are significantly above all theory curves, except the mean-field result; no sign of decrease toward the BCS regime can be discerned. This suggests that some other factors, e.g., nonequilibrium, finite size effect, need to be considered. Therefore, we shall refrain from comparing with these experimental data in the BCS regime. In short, our results with the particle-hole contributions are consistent with experiment in the unitary and BEC regimes and in good agreement with the QMC results [23] in the unitary and BCS regimes. This calls for more data from future experiment.
A few remarks are in order. (i) In the BCS limit, the bosonic fluctuation contributions are relatively weak. The mean-field superfluid density $n_{s}/m$ has a strong temperature dependence and decreases sharply toward zero at the mean-field transition temperature $T_{c}^{\text{MF}}$. Any fluctuation effect would only suppress $n_{s}/m$. Therefore, it serves to provide an upper bound for $T_{\text{BKT}}$. The sharp $T$ dependence of $n_{s}/m$ (and $n_{\text{B}}/M_{\text{B}}$) near $T_{c}^{\text{MF}}$ implies that one has $T_{\text{BKT}}\lesssim T_{c}^{\text{MF}}$, insensitive to the actual condition used. (ii) It is well known that, in the continuum, one has $n_{s}(T=0)/m=n/m$ due to the Galilean translational invariance. For this reason, Shi et al. [46] presented an interaction-independent upper bound $T_{F}/8$ for $T_{\text{BKT}}$. Thus it is hard to imagine that $n_{s}(T=T_{\text{BKT}})/m$ would properly reflect the effect of inter-fermion interactions. In this sense, we suspect that the NKC cannot properly capture the interaction effect in the crossover regime of a fermionic superfluid system. (iii) Furthermore, in an NKC-based approach [4, 37], it is often assumed $M_{B}=2m$ throughout the BCS-BEC crossover, ignoring possible pair mass renormalization. (iv) Therefore, the non-monotonic behavior of $T_{\text{BKT}}$ we find in Fig. 4(a) is unlikely to be present in an NKC-based calculation.
@@
Finally, we note that the particle-hole susceptibility in Eq. (9) is density independent in 2D. This suggests that the particle-hole contributions in the weak coupling limit would have already been automatically included, if the 2D scattering length $a_{\text{2D}}$ were measured directly through experiment, e.g., by measuring the two-body binding energy in the dilute limit. Instead, $a_{\text{2D}}$ is usually calculated from the 3D scattering length as in a deeply confined pancake-shaped trap [40], thus the effect of particle-hole fluctuations should be seriously taken into account when comparing experiment and theory. In Fig. 6, the same definition of $a_{\text{2D}}$ was used in both QMC and our present work. This makes it possible to have a good agreement in the BCS regime between QMC and our result with particle-hole fluctuations included. Conversely, one can experimentally determine the particle-hole susceptibility in the BCS limit by measuring the 2D scattering length and comparing with that calculated from $a_{\text{3D}}$, and obtain
@@
where $a^{\text{exp}}_{\text{2D}}$ denotes the experimentally measured 2D scattering length. Nevertheless, away from the weak coupling limit, a nontrivial density dependence should emerge as both a nonzero pairing gap develops and the ratio $\mu/E_{\text{F}}$ starts to decrease with increasing pairing strength at a finite density.
@@ IV Conclusions
In summary, we have studied the impact of the particle-hole channel on the BKT physics in Fermi gases within the context of BCS-BEC crossover. The entire particle-hole $T$ matrix is included to renormalize the interaction that appears in the particle-particle $T$ matrix. Both $T$ matrices contain self-consistently includes the self-energy feedback. This leads to a non-uniform shift of the inverse interaction strength by the particle-hole susceptibility, which evaluated as an angular average at two different levels, revealing important physical consequences in the crossover and BCS regimes. The particle-hole channel provides a screening of the pairing interaction and thus shifts curves of the BKT transition temperature and the pairing gap towards the BEC regime as a function of $\ln(k_{F}a_{\text{2D}})$. Additionally, a comparison shows that the particle-hole corrected BKT transition temperature exhibits a better agreement with the experimental data and QMC results. Future experiments with more elaborate measurements are called for to resolve various discrepancies.
Finally, it should be mentioned that, in addition to the simple averaged particle-hole susceptibility, there are a series of higher order corrections, including a higher order $T$-matrix [17] and vertex corrections, which may contribute significant modifications to the present results. In addition, the BKT criterion [50, 49] is merely based on numerical simulations [41] to provide the critical phase space density. Despite the experimental support [38], a rigorous analytical derivation of such a criterion would be highly desirable.
@@ Appendix A Slope discontinuity in $\chi_{\text{ph}}$ across $\mu=0$ at zero $T$
Appendix B Determining the critical phase space density ${\cal D}_{B}$
Estimates of $\mathcal{D}_{\text{B}}(T_{\text{BKT}})$ for fermionic superfluids are provided in Refs. [44, 38], with values ranging from approximately 4.9 to 6.45. In comparison, the analogous systems in atomic Bose gases typically hover around 8 [48]. The lowest value, $\mathcal{D}_{\text{B}}(T_{\text{BKT}})=4.9$, which is the closest to the factor of 4 in the usual Nelson-Kosterlitz BKT relation, was reported to offer the best fit for the experimental results on Fermi gases [38]. In Ref. [49], $\mathcal{D}_{\text{B}}(T_{\text{BKT}})$ was taken at fixed value of 4.9. Here we consider a broader range of interaction and thus include the loglog dependence, albeit very weak, of $\mathcal{D}_{\text{B}}$ on $a_{\text{2D}}$, and take the data from experiment. To this end, we take the data of $\tilde{g}$ from the Supplemental Materials of Ref. [38], and plot in Fig. 7. The data fits perfectly to Eq. (6) in the text. Also plotted in the figure is the function with the identification of potential radius [41] or pair size [39] with $a_{\text{2D}}$. The small difference between these two curves reveals that $a_{\text{2D}}$ differs from either potential radius or pair size.
Figure 7: $\tilde{g}$ as a function of $\ln(k_{F}a_{\text{2D}})$ as given by the QMC result of Prokof’ev et al [41] (black solid line). Shown in the inset is the corresponding ${\cal D}_{\text{B}}$ (red line), along with the experimental data (green dots) and fit (green line) for $\tilde{g}$ from Ref. [38]. The fit from experimental data yields a slightly larger ${\cal D}_{\text{B}}$ (blue line).